OpenStudy (anonymous):

consider the following quadratic equation: 9x^2+6x+1=0 using the standard form ax^2+bx+c=0 of the quadratic equation above, factor the left hand side of the equation into two linear factors.

myininaya (myininaya):

9x^2+6x+1=0 9x^2+3x+3x+1=0 3x(3x+1)+1(3x+1)=0 (3x+1)^2=0

OpenStudy (anonymous):

(3x+1)^2=0